OpenStudy (anonymous):

WILL REWARD A beam rests against a wall, forming a 65 degree with the floor. Use the function y=9sec theta to find the length of the beam to the nearest tenth of a foot. A. 8.2 ft B. 21.3 ft C. 9.9 ft D. 18.7 ft

OpenStudy (anonymous):

How secant relates to the sides of a right triangle: \[Sec(\theta) = \frac{1}{Cos(\theta)}= \frac{Hypotenuse}{Adjacent}\] Therefore: Hypotenuse = Adjacent * sec( theta ) When we draw a picture, we see that the angle given to us is the correct angle. So, we simply put the 65 degree angle in for theta, and get our answer out :) What is 9 * sec(65)?

OpenStudy (anonymous):

I got 10.986

OpenStudy (anonymous):

@Vandreigan

OpenStudy (anonymous):

That's not quite what I get.

OpenStudy (anonymous):

Shouldn't be 9cos(65)?

OpenStudy (anonymous):

@Vandreigan

OpenStudy (anonymous):

Not cos. We don't know the length of the ladder, as I'm understanding the question. That's what we want to find out.

OpenStudy (anonymous):

Okay i got 21.29 for 9*sec(65)

OpenStudy (anonymous):

 That's the setup of the question, as I'm reading it :)